Consider the linear transformation T: P2 R3 given by . T(p) = p(0) p(-3) p(2) 3 Solve T(p) = 9 Show your calculations.
Answer #1

solution : 0 given that The lincat Townsformation T: 13-> (PLO) T(P) - P(-3) PLO) let pe - ast batc - 6 PCO)- a (0) + 6(0)+C30 C value in Equalia @ © qa-3b+3=9 qa-36=9-3 qa-36=6 2 ilja-b) - 6 za-b=2 č value in Equation ③ 40+26+3=-1 4a+26=-1-3 ua+

I HOPE THIS ANSWER WILL BE HELPFUL TO YOU

PLEASE LIKE THIS ANSWER

THANK YOU.